Consider sending a datagram of size 2400 bytes into a link whose MTU is 700. How many fragmentations are generated? Show the length, fragflag and offset fields of each IP datagram after fragmentation.
Answer #1

Dafagram of size=2400 bytes HTo = Too byko но Феотипи = 120 | С( IP header fire = 20 bytes needs to be subtracred ) = 73.57-4 so 4 fragments are needed each fragment size= 680 bytes (s: Itsince 680 +680 +680+ 340 = 2380 - t 3 fragmets 4th fregunt (e) Dafagram sie = 2380 ip) 2400- 20 - 2380 Ip headee ME (More fra